# Identifying Equivalent Training Dynamics

### v2

@@ D.2 Computing Koopman mode decomposition
We found in Sec. 4.4 that the Transformers trained with constrained weight norm (and do not undergo grokking) have non-conjugate dynamics with the Transformers trained with unconstrained weight norm (and do undergo grokking). To ensure that this was not due to the fact that training loss over the window of training time used for computing the Koopman eigenvalues (the first $T=100$ training iterations) was distinct between the two Transformers, we performed the following control experiment. Namely, we performed the same analysis, but we considered the weight trajectories of Transformers with constrained weight norm over twice as long a training time interval ($T=200$) In this case, the constrained Transformer reaches a training loss that is closer to that of the unconstrained Transformer (Fig. S3A – compare solid red and black lines), although there is more variability between independent seeds (Fig. S3A – shaded red area). However, in this case we again find that the Koopman eigenvalues associated with unconstrained and constrained training are distinct (Fig. S3B). In particular, the constrained Transformer again has a pair of complex conjugate Koopman eigenvalues that lie along the unit circle, while the unconstrained Transformer has a pair of complex conjugate Koopman eigenvalues outside of the unit circle. This suggests distinct stability properties, and further emphasizes the lack of conjugacy.
@@
Figure S3: Distinct Koopman eigenvalues between Transformers that do and do not undergo grokking is consistent, when considering windows of training time that have more similar training loss. (A) As in Fig. 6C, we train Transformers with unconstrained weight norm for $T=100$ (thin vertical black line) iterations and use the associated weight trajectories to approximate the KMD. In contrast, we train Transformers with constrained weight norm for $T=200$ iterations (thin vertical red line) and use the associated weight trajectories to approximate the KMD. This enables the two networks to reach more comparable (although not perfectly matching) training loss. Lines are mean across $20$ indepedently trained Transformers and shaded area is $\pm$ one standard deviation. (B) Same as Fig. 6D, but with the Koopman eigenvalues associated with training the constrained Transformer for $T=200$ iterations.

### v3

@@ E.2 Computing Koopman mode decomposition
We found in Sec. 4.4 that the Transformers trained with constrained weight norm (that do not undergo grokking) have non-conjugate dynamics with the Transformers trained with unconstrained weight norm (that do undergo grokking). To ensure that this was not due to the fact that training loss over the window of training time used for computing the Koopman eigenvalues (the first $T=100$ training iterations) was distinct between the two Transformers, we performed the following control experiment. Namely, we performed the same analysis, but we considered the weight trajectories of Transformers with constrained weight norm over twice as long a training time interval ($T=200$). In this case, the constrained Transformer reaches a training loss that is closer to that of the unconstrained Transformer (Fig. S6A – compare solid red and black lines), although there is more variability between independent seeds (Fig. S6A – shaded red area). However, in this case we again find that the Koopman eigenvalues associated with unconstrained and constrained training are distinct (Fig. S6B). In particular, the constrained Transformer again has a pair of complex conjugate Koopman eigenvalues that lie along the unit circle, while the unconstrained Transformer has a pair of complex conjugate Koopman eigenvalues outside of the unit circle. This suggests distinct stability properties, and further emphasizes the absence of a conjugacy.
@@
Figure S6: Distinct Koopman eigenvalues between Transformers that do and do not undergo grokking is consistent when considering windows of training time that have more similar training loss. (A) As in Fig. 5C, we train Transformers with unconstrained weight norm for $T=100$ (thin vertical black line) iterations and use the associated weight trajectories to approximate the KMD. In contrast, here we train Transformers with constrained weight norm for $T=200$ iterations (thin vertical red line) and use the associated weight trajectories to approximate the KMD. This enables the two networks to reach more comparable (although not perfectly matching) training losses. Lines are mean across $20$ independently trained Transformers and shaded area is $\pm$ one standard deviation. (B) Same as Fig. 5D, but with the Koopman eigenvalues associated with training the constrained Transformer for $T=200$ iterations.